# ESP-MedSAM: Efficient Self-Prompting SAM for Universal Domain-Generalized Medical Image Segmentation

### v4

@@ 3.3 Self-Patch Prompt Generator
Current medical SAM [14, 15, 17, 19] and lightweight SAM [33, 37, 38, 36, 35, 39] mainly leverage manual prompts (e.g., points and boxes) to guide the model providing satisfied segmentation masks. However, such methods rely on the experience of pathologists in medical scenarios, which are expensive and time-consuming. To eliminate the demand for manual annotations, we propose the SPPG that contains a patch generator and Dense Prompt (DP) encoder to automatically produce a set of high-quality patch prompts to assist segmentation decoding. As presented in Fig. 2, the patch generator $\delta$ consists of $L$ convolutional layers for patch merging and $1\times 1$ convolution for channel compression, where each layer includes $2\times 2$ convolution with stride 2 followed by LayerNorm. With each additional convolutional layer, the scale of the patch is doubled. The prediction of patch prompts $\hat{\mathcal{P}}$ can be formulated as:
$\hat{\mathcal{P}}=\frac{1}{1+e^{(-\delta(\textit{h}))}},$ (5)
where $\hat{\mathcal{P}}\in\mathbb{R}^{\frac{H\times W/L}{m}}$. As the image embedding h contains rich semantic information and the target area of many medical images is much smaller than the background, we place the patch generator on h. The patch prompt is essentially an inductive prediction, reducing the complexity of predicting the final segmentation mask by providing the summarized semantic information to the decoder. To optimize the generated patch prompts, we utilise binary cross-entropy loss to perform computation:
$L_{\rm SPPG}=-\frac{1}{K}\sum_{k=1}^{K}\mathcal{P}^{k}\log(\hat{\mathcal{P}}^{k})+(1-\mathcal{P}^{k})\log(1-\hat{\mathcal{P}}^{k}),$ (6)
@@
where $\mathcal{P}^{k}$ is the target patch generated by maxpooling the ground truth of source domains $\mathcal{S}^{k}$. We then respectively use the interpolation method for upsampling $\hat{\mathcal{P}}_{k}$ and $1\times 1$ convolution to align the dimension with the image embedding h. They constitute the DP encoder. In this way, the SPPG module automatically produces a set of high-quality dense prompts p to guide the prediction of segmentation masks, improving the applicability of ESP-MedSAM in clinical scenarios.
Table 1: Details of the source domains used in our experiments.
@@
$\mathcal{S}^{2}$ PCXA [42, 43] X-ray Variable 704
@@
Table 2: Details of the unseen domains used in our experiments.
@@
Table 3: Comparison with State-of-the-art Frameworks in Universal Medical Image Segmentation (source domains).
@@
ESP-MedSAM 88.52 92.42 96.83 40.68 82.42 62.64 92.93 56.32 85.28 82.32 92.24 85.93
Table 4: Comparison of Computation Costs with Automatic Segmentation Models under 1024 $\times$ 1024 Inputs.
Method Params $\downarrow$ FLOPs $\downarrow$ FPS $\uparrow$
U-Net [3] 13.40M 497.91G 11.07
U-Net++ [18] 9.16M 558.46G 8.57
Att-UNet [4] 8.72M 269.07G 8.45
nnUNet [5] 34.29M 554.40G 10.99
H2Former [6] 33.71M 536.96G 7.41
TransUNet [54] 109.54M 865.31G 4.73
ADS [9] 14.07M 503.59G 10.91
CIAug [8] 14.26M 500.23G 10.85
MADGNet [7] 31.40M 210.65G 8.33
MobileSAM [33] 9.79M 39.71G 1.05
RepViT-SAM [37] 9.57M 23.64G 1.12
EfficientViT-SAM [38] 34.80M 89.10G 0.91
EfficientSAM [36] 25.38M 32.51G 0.98
EdgeSAM [35] 9.60M 22.10G 1.25
SAM-Lightening [39] 19.26M 52.46G 1.02
ESP-MedSAM 28.46M 55.88G 12.87
3.4 Query-Decoupled Modality Decoder
[…]
$\mathcal{H}=\mathrm{softmax}(\frac{(\textit{h'}\oplus\Omega)\cdot(q_{\rm sa}^{k})^{T}}{\sqrt{d}})\cdot q_{\rm sa}^{k}\oplus\textit{h'},$’ , (7)
where $\mathcal{H}$ is the decoding embedding, $\Omega$ stands for the corresponding positional encodings. Similar to SAM [13], such operations are iterated twice. To predict the segmentation mask $\hat{y}^{k}_{i}$, we perform:
$\hat{y}^{k}_{i}=\sigma(\mathcal{F}_{\rm conv}(\mathcal{H})\cdot\Phi^{k}_{\rm dec}(q_{\rm sa}^{k})),$ (8)
where $\mathcal{F}_{\rm conv}$ represents two $2\times 2$ transpose convolutions, upsampling $\mathcal{H}$ by 2$\times$. $k$-th MLP $\Phi^{k}_{\rm dec}$ aligns the channel with upscaled $\mathcal{H}$ and contains the decoding information of corresponding modalities. $\sigma$ stands for a sigmoid operation followed by an interpolation function that recovers the size of masks. Consequently, the proposed QDMD provides an independent decoding process for each modality, which avoids the conflicting inherent heterogeneity of different modalities, improving the generalization capabilities of our ESP-MedSAM framework.
Table 5: Comparison with State-of-the-art Frameworks in Domain-Generalized Medical Image Segmentation (unseen domains).
@@ 3.4 Query-Decoupled Modality Decoder
ESP-MedSAM 91.45 78.13 92.24 91.30 79.68 83.84 65.96 160.13 61.62 199.74 64.21 248.47
4 Experiments
4.1 Datasets and Implementations
4.1.1 Datasets
To validate the effectiveness of our proposed ESP-MedSAM, we select six different medical imaging modalities: dermoscopy, X-ray, fundus, colonoscopy, ultrasound and microscopy. Table 1 presents the datasets of source domains for proving the universal ability. To evaluate the domain generalizability of our framework, we further collect an external dataset for each modality, forming a set of unseen domains, as illustrated in Table 2. We follow official guidelines for all datasets to create our training, validation and test sets.
Table 6: Ablation study of ESP-MedSAM in domain-generalized Medical Image Segmentation: $\mathcal{S}\rightarrow\mathcal{T}$.
Row MMDKD SPPG QDMD Dice (Avg.) $\uparrow$ HD (Avg.) $\downarrow$ Params $\downarrow$
1 64.49 451.43 90.58M
2 73.81 302.72 29.08M
3 69.96 152.65 90.60M
4 67.28 297.51 89.94M
5 74.29 142.97 29.10M
6 75.17 272.88 28.44M
7 70.83 143.26 89.96M
8 75.86 140.60 28.46M
4.1.2 Implementation Details
We perform all experiments on a single NVIDIA A6000 GPU with PyTorch. We adopt the optimizer using Adam with a learning rate of $1\times 10^{-4}$. The batch size and epochs are set to 4 and 200, respectively. We apply the exponential decay strategy to adjust the learning rate, where the factor is set as 0.98. All images are resized to $1024\times 1024$ during the training and test stages. We set the patch size $m$ to be $32$, $L=1$ and utilize 1% images of the SA-1B dataset for the distillation. We use ViT-H [12] as the teacher image encoder for all SAM-based architectures. The loss coefficient $\lambda$ is set to 0.7 during training. In comparing traditional automatic segmentation methods, we consider both UNet-based [3, 18, 4, 5, 6, 54] and DG-based [8, 9, 7] fully-automated architectures as baselines. In the comparison of lightweight SAM frameworks, MobileSAM [33], RepViT-SAM [37], EfficientViT-SAM [38], EfficientSAM-Ti [36], EdgeSAM [35] and SAM-Lightening [39] are served as baselines. These architectures are trained with the point prompt mode that uses the ConnectedComponentsWithStats function in OpenCV to calculate the centroid of each instance as point prompts. In addition to evaluating the manual point mode, they perform the automatic mask generation mode [13].
4.1.3 Evaluation Metrics
To perform the comprehensive evaluation of medical image segmentation, we adopt two common metrics: Dice coefficient and Hausdorff Distance (HD). Both measure the similarity between the prediction and Grouth Truth (GT), where HD is more sensitive to the boundary than Dice. We also report model parameters, Floating Point Operations (FLOPs) and Frames Per Second (FPS) to reveal the computation costs. To match the predicted $K$ masks with the corresponding ground truth $y^{k}_{i}$, we calculate the Dice score between $\{\hat{y}^{k}_{i}\}_{k=1}^{K}$ and $y^{k}_{i}$. The one with the highest Dice score in this set is chosen as the matching prediction mask for the following evaluation of segmentation [14].
4.2 Comparison on Universal Medical Image Segmentation
@@
Figure 3: Qualitative comparison with SOTA task-specific models and lightweight SAM frameworks with the automatic mask generation mode on medical image segmentation over source and unseen target domains of six modalities.
To validate the effectiveness of our lightweight ESP-MedSAM framework in universal medical instance segmentation, we compare it with SOTA fully-automated architectures: U-Net [3], U-Net++ [18], Att-UNet [4], nnUNet [5], H2Former [6], TransUNet [54], ADS [9], CIAug [8], MADGNet [7] and lightweight SAMs [33, 37, 38, 36, 35, 39] on source domains of six medical imaging modalities, as illustrated in Table 3. For fair comparisons, the knowledge of lightweight SAMs is distilled from the same teacher model $\mathcal{F}_{\rm ViT-Med}$. Firstly, it can be observed that ESP-MedSAM performs better than all task-specific models, especially achieving a Dice of 85.28% in $S^{5}$ and being 10.06% higher than nnUNet [5]. Note that all task-specific models require to be retrained on each domain. Therefore, our ESP-MedSAM demonstrates outstanding universal capabilities. Secondly, compared to recent lightweight SAMs, the proposed ESP-MedSAM outperforms the laborious point-prompt mode. Particularly, with the automatic mask generation mode, the performance of these lightweight SAMs in retinal vessel and nuclei segmentation tasks declines rapidly. The proposed ESP-MedSAM significantly surpasses EfficientViT-SAM [38] with a 56.47% and 25.92% Dice increase, respectively. On HD metric, our framework reduces the distance by up to 5.8$\times$ compared to current SOTA models, revealing more accurate boundary prediction. Furthermore, we present the computation costs of each framework in Table 4. In high-resolution medical image segmentation tasks, we observe that traditional task-specific architectures demonstrate acceptable inference speed but higher model complexity. The recent lightweight SAMs suffer from the slow inference speed in the automatic mask generation mode due to the inefficient sliding window algorithm. In contrast, our ESP-MedSAM displays remarkable complexity-speed trade-offs. These results reveal the superior universal capabilities of our ESP-MedSAM on diverse medical segmentation tasks without the demand for manual prompts.
@@
4.3 Comparison on Domain-Generalized Medical Image Segmentation

### v5

@@ III-C Self-Patch Prompt Generator
Current medical SAM [17, 18, 4, 28] and lightweight SAM [22, 38, 24, 23, 37, 39] mainly leverage manual prompts (e.g., points and boxes) to guide the model to provide satisfactory segmentation masks. However, such methods rely on the experience of pathologists in medical scenarios, which are expensive and time-consuming. To eliminate the demand for manual annotations, we propose the SP-Generator that contains a patch generator and prompt encoder to automatically produce a set of high-quality patch prompts to assist segmentation decoding. Specifically, the patch generator $\delta$ consists of $L_{2}$ convolutional layers for patch merging and $1\times 1$ convolution $\mathcal{F}_{\rm Conv}^{\rm 1\times 1}$ for channel compression, where each layer includes $2\times 2$ convolution $\mathcal{F}_{\rm Conv}^{\rm 2\times 2}$ with stride 2 followed by LayerNorm. With each additional convolutional layer, the scale of the patch is doubled. After that, we utilize the sigmoid $\sigma$ function to generate probability maps. The self-predicted patch prompts $\hat{p}_{\rm pat}$ is optimized by binary cross-entropy loss since this prediction can be considered as a global classification task:
$L_{\rm pat}=-p_{\rm pat}\log(\hat{p}_{\rm pat})+(1-p_{\rm pat})\log(1-\hat{p}_{\rm pat}),$ (3)
@@
where $p_{\rm pat}$ is the target patch generated by the maxpooled ground truth. We then respectively use the interpolation method for memory-efficient upsampling $\hat{p}_{\rm pat}$ and $1\times 1$ convolution to align the dimension with the image embedding $h$. They constitute the prompt encoder. During inference, the computation of dense prompt embeddings $p_{\rm dense}$ can be formulated as:
$p_{\rm dense}=\mathcal{F}_{\rm Conv}^{\rm 1\times 1}(\mathcal{F}_{\rm inter}(\sigma(\mathcal{F}_{\rm Conv}^{\rm 1\times 1}(\mathcal{F}_{\rm LN}(\mathcal{F}_{\rm Conv}^{\rm 2\times 2}(h))))))).$ (4)
As the image embedding h contains rich semantic information and the target area of many medical images is much smaller than the background, we place the patch generator on h. The patch prompt is essentially an inductive prediction, reducing the complexity of predicting the final segmentation mask by providing the summarized semantic information to the decoder. In this way, the SP-Generator module automatically produces a set of high-quality dense prompts to guide the prediction of segmentation masks, improving the applicability of De-LightSAM in clinical scenarios.
III-D Query-Decoupled Modality Decoder
The mask decoder of SAM [16] utilizes modality-agnostic query tokens to handle all segmentation tasks in natural images. However, this is not optimal for medical image segmentation. As there exists inherent heterogeneity in various medical imaging modalities, such common prediction channels suffer from decoding conflicts, degrading the generalization capabilities of the model. To address this problem, we propose the QM-Decoder for our De-LightSAM framework. Concretely, we set $K$ query tokens $\{q^{k}\}^{K}_{k=1}$, where $q^{k}\in\mathbb{R}^{1\times C_{2}}$, to customize the private workflow for each Modality. We adopt a class-fixed assignment algorithm where each query token corresponds to one modality category. Given the image embeddings h and prompt embeddings p, we first update the mask query using the self-attention operation and then conduct cross-attention with the fusion of h and $p_{\rm dense}$: $h\leftarrow h\oplus p_{\rm dense}$, where $\oplus$ is an element-wise addition operation:
$q^{k}\leftarrow\mathrm{softmax}(\frac{\mathcal{Q}(q^{k})\cdot\mathcal{K}(q^{k})^{T}}{\sqrt{d}})\cdot\mathcal{V}(q^{k}),$ (5)
$h^{\prime}=\mathrm{softmax}(\frac{\mathcal{Q}(h\oplus\Omega)\cdot\mathcal{K}(q^{k})^{T}}{\sqrt{d}})\cdot\mathcal{V}(h)\oplus h,$ (6)
where $h^{\prime}$ is the decoding embedding, $\Omega$ stands for the corresponding positional encodings. Similar to SAM [16], such operations are iterated twice. To predict the segmentation mask $\hat{y}^{k}$, we perform:
$\hat{y}^{k}=\mathcal{F}_{\rm inter}(\sigma(\mathcal{F}^{2\times 2}_{\rm trans}(\mathcal{F}^{2\times 2}_{\rm trans}(h^{\prime}))\cdot\mathcal{F}^{k}_{\rm MLP}(h^{\prime})),$ (7)
where $\mathcal{F}^{2\times 2}_{\rm trans}$ represents two $2\times 2$ transpose convolutions, upsampling $h^{\prime}$ by 4$\times$. $k$-th MLP $\mathcal{F}^{k}_{\rm MLP}$ aligns the channel with upscaled $\mathcal{H}$ and contains the decoding information of corresponding modalities. The sigmoid operation, followed by an interpolation function, is used to recover the size of the original masks. The quality of predicted masks is supervised by a combination of focal loss and dice loss:
$L_{\rm seg}=L_{\rm focal}+L_{\rm dice}.$ (8)
Consequently, the proposed QM-Decoder provides an independent decoding process for each modality, which avoids the conflicting inherent heterogeneity of different modalities, improving the generalization capabilities of our De-LightSAM.
III-E Multi-Modal Decoupled Knowledge Distillation
Knowledge distillation has emerged as a prominent technique for transferring knowledge from large foundation models to lightweight architectures while maintaining competitive performance. To compress the SAM size, recent studies [37, 23, 22] usually adopt the same modality as the pretraining during the distillation stage. However, the scarcity of medical imaging data poses additional challenges for effective knowledge distillation. Medical datasets are inherently limited in scale due to privacy concerns, expensive annotation costs, and the specialized expertise required for accurate labeling. This data scarcity leads to insufficient sample diversity for robust feature learning, causing the student model to overfit to the limited training samples and fail to generalize to unseen medical domains. To address this issue, we propose an MDKD strategy to provide robust foundational knowledge that can complement domain-specific medical features. Specifically, we construct a multi-modal teacher model that involves two complementary components: (1) a natural image encoder $\mathcal{F}_{\rm{ViT\text{-}SAM}}$ of SAM [16] to provide common knowledge as this image encoder is pretrained on a large-scale dataset containing a wide variety of diversity and (2) a medical SAM image encoder $\mathcal{F}_{\rm{ViT\text{-}Med}}$ to provide domain-specific knowledge. To obtain $\mathcal{F}_{\rm{ViT\text{-}Med}}$, we integrate an adapter-based SAM encoder with our SP-Generator and QM-Decoder, adopting the parameter-efficient fine-tuning technique [41] to achieve adaptation on source domains. In particular, to reduce the number of learnable parameters, we only insert adapters in FFN layers as they determine the feature generation. Drawing inspiration from the divide-and-conquer algorithm, our proposed MDKD method decouples the KD process into two sub-tasks: common feature distillation and domain-specific feature distillation. During common feature distillation, we aim to transfer general visual representation knowledge from $\mathcal{F}_{\rm{ViT\text{-}SAM}}$ to the backbone components of DC-Encoder (denoted as $\mathcal{F}_{\rm{DCE}}$) without involving the modality controller. This process ensures that our lightweight encoder can capture fundamental visual patterns and structural information that are universally applicable across different imaging modalities. Moreover, we distill the specialized knowledge of each medical modality from $\mathcal{F}_{\rm{ViT\text{-}Med}}$ to the modality controller $\mathcal{F}_{\rm{DCE}}^{\rm MC}$ of our DC-Encoder. This targeted distillation enables the modality controller to learn domain-specific characteristics and adapt the shared backbone features to the unique requirements of each medical imaging modality. By separating this process from common feature distillation, we ensure that modality-specific adaptations do not interfere with the general visual representation capabilities. The loss function of MDKD is defined by:
$L_{\rm MDKD}$ $=\big{|}\big{|}\mathcal{F}_{\rm{ViT\text{-}SAM}}(\textbf{{x}}_{\rm nat})-\mathcal{F}_{\rm{DCE}}(\textbf{{x}}_{\rm nat})\big{|}\big{|}^{2}_{2}$ (9)
$+\sum^{K}_{k=1}\big{|}\big{|}\mathcal{F}_{\rm{ViT\text{-}Med}}(\textbf{{x}}^{k})-\mathcal{F}^{\rm MC}_{\rm{DCE}}(\textbf{{x}}^{k})\big{|}\big{|}^{2}_{2},$
where $\textbf{{x}}_{\rm nat}$ stands for the $1\%$ natural images of the SA-1B dataset [16] and $\textbf{{x}}^{k}$ is the medical image collected from source domains $\mathcal{S}^{k}$. Overall, our MDKD method incorporates large-scale natural image data to provide robust common knowledge, leading to superior segmentation performance and enhanced domain generalization capabilities.
TABLE I: Details of the source domains used in our experiments.
@@ III-E Multi-Modal Decoupled Knowledge Distillation
$\mathcal{S}^{2}$ PCXA [44] X-ray Variable 704
@@
TABLE II: Details of the unseen domains used in our experiments.
@@
TABLE III: Comparison with State-of-the-art Frameworks in Universal Medical Image Segmentation (source domains).
@@
De-LightSAM 88.52 92.42 96.83 40.68 82.42 62.64 92.93 56.32 85.28 82.32 92.24 85.93
Figure 4: Qualitative comparison with SOTA task-specific models and lightweight SAM frameworks on medical image segmentation (source domain).
TABLE IV: Comparison with State-of-the-art Frameworks in Domain-Generalized Medical Image Segmentation (unseen domains).
@@
De-LightSAM 91.45 78.13 92.24 91.30 79.68 83.84 65.96 160.13 61.62 199.74 64.21 248.47
@@
Figure 5: Qualitative comparison with SOTA task-specific models and lightweight SAM frameworks on medical image segmentation (target domain).
TABLE V: Comparison of Computation Costs with Automatic Segmentation Models under 1024 $\times$ 1024 Inputs.
Method Params $\downarrow$ FLOPs $\downarrow$ FPS $\uparrow$ Latency $\downarrow$
U-Net [5] 13.40M 497.91G 11.07 90.34ms
U-Net++ [25] 9.16M 558.46G 8.57 116.69ms
Att-UNet [6] 8.72M 269.07G 8.45 118.34ms
nnUNet [7] 34.29M 554.40G 10.99 91.00ms
H2Former [8] 33.71M 536.96G 7.41 134.95ms
TransUNet [55] 109.54M 865.31G 4.73 211.42ms
ADS [14] 14.07M 503.59G 10.91 91.67ms
CIAug [13] 14.26M 500.23G 10.85 92.17ms
MADGNet [12] 31.40M 210.65G 8.33 120.05ms
MobileSAM [22] 9.79M 39.71G 1.05 952.38ms
RepViT-SAM [38] 23.16M 115.61G 1.12 892.86ms
EfficientViT-SAM [24] 34.80M 89.10G 0.91 1098.90ms
EfficientSAM [23] 25.38M 32.51G 0.98 1020.41ms
EdgeSAM [37] 9.60M 22.10G 1.25 800.00ms
SAM-Lightening [39] 19.26M 52.46G 1.02 980.39ms
De-LightSAM 12.74M 55.86G 13.09 76.39ms
@@
TABLE VI: Ablation study of De-LightSAM in domain-generalized Medical Image Segmentation: $\mathcal{S}\rightarrow\mathcal{T}$.
Row $M_{1}$ $M_{2}$ $M_{3}$ Dice (Avg.) $\uparrow$ HD (Avg.) $\downarrow$ Params $\downarrow$
1 64.49 451.43 641.09M
2 71.94 286.32 13.36M
3 69.96 152.65 641.11M
4 67.28 297.51 640.45M
5 74.29 142.97 13.38M
6 75.17 272.88 12.72M
7 70.83 143.26 640.47M
8 75.86 140.60 12.74M
TABLE VII: Ablation study of MDKD in domain-generalized Medical Image Segmentation: $\mathcal{S}\rightarrow\mathcal{T}$.
Row Method Dice (Avg.) $\uparrow$ HD (Avg.) $\downarrow$
1 DC-Encoder 66.17 192.49
2 +KD 72.04 165.23
3 +MDKD 75.86 140.60
IV Experiments
IV-A Datasets and Implementations
IV-A1 Datasets
To validate the effectiveness of our proposed De-LightSAM, we select six different medical imaging modalities. Table I presents the datasets of source domains for proving the universal ability. We adopt a common train-val-test split of 8:1:1 in PCXA, CVC-ClinicDB, UDIAT and DSB-2018. For ISIC-2018  [^1] https://challenge.isic-archive.com/data/#2018 and DRIVE  [^2] https://drive.grand-challenge.org/ datasets, we use the official train-val-test sets provided by their competition organizers. To evaluate the generalization capability of our framework, we further collect an unseen domain for each modality, as illustrated in Table II.
IV-A2 Implementation Details
We perform all experiments on a single NVIDIA A6000 GPU with PyTorch. We adopt the optimizer using Adam with a learning rate of $1\times 10^{-4}$. The batch size and epochs are set to 4 and 200, respectively. We apply the exponential decay strategy to adjust the learning rate, where the factor is set as 0.98. All images are resized to $1024\times 1024$ during the training and test stages. We set the patch size $m$ to be $32$, $L_{2}=1$ and utilize 1% images of the SA-1B dataset for the distillation. We use ViT-H [15] as the teacher image encoder for all SAM-based architectures. The loss coefficient $\lambda$ is set to 0.7 during training. In comparing traditional automatic segmentation methods, we consider both UNet-based [5, 25, 6, 7, 8, 55] and DG-based [13, 14, 12] fully-automated architectures as baselines. We reimplement these models, follow their official configurations in our experiments, and train a single model for each domain. In the comparison of lightweight SAM frameworks, MobileSAM [22], RepViT-SAM [38], EfficientViT-SAM [24], EfficientSAM-Ti [23], EdgeSAM [37] and SAM-Lightening [39] are served as baselines. These architectures are trained with the point prompt mode that uses the ConnectedComponentsWithStats function in OpenCV to calculate the centroid of each instance as point prompts. In addition to evaluating the manual point mode, they perform the automatic mask generation mode [16].
IV-A3 Evaluation Metrics
To perform the comprehensive evaluation of medical image segmentation, we adopt two common metrics: Dice coefficient and Hausdorff Distance (HD). Both measure the similarity between the prediction and ground truth, where HD is more sensitive to the boundary than Dice. We also report model parameters, Floating Point Operations (FLOPs), Frames Per Second (FPS) and latency to reveal the computation cost and inference speed. To match the predicted $K$ masks with the corresponding ground truth $y^{k}_{i}$, we calculate the Dice score between $\{\hat{y}^{k}_{i}\}_{k=1}^{K}$ and $y^{k}_{i}$. The one with the highest Dice score in this set is chosen as the matching prediction mask for the evaluation [17].
IV-B Comparison with State-of-the-arts on Source Domains
To validate the effectiveness of our lightweight De-LightSAM framework in universal medical segmentation, we compare it with SOTA fully-automated architectures: U-Net [5], U-Net++ [25], Att-UNet [6], nnUNet [7], H2Former [8], TransUNet [55], ADS [14], CIAug [13], MADGNet [12] and lightweight SAMs [22, 38, 24, 23, 37, 39] on the test set of source domains, as illustrated in Table III. For fair comparisons, the knowledge of lightweight SAMs is distilled from the same teacher model $\mathcal{F}_{\rm ViT-Med}$. Firstly, it can be observed that De-LightSAM performs better than all task-specific models, especially achieving a Dice of 85.28% in $S^{5}$ and being 10.06% higher than nnUNet [7]. Note that all task-specific models require to be retrained on each domain. Therefore, our De-LightSAM demonstrates outstanding universal capabilities. Secondly, compared to recent lightweight SAMs, the proposed De-LightSAM outperforms the laborious point-prompt mode. Particularly, with the automatic mask generation mode, the performance of these lightweight SAMs in retinal vessel and nuclei segmentation tasks declines rapidly. The proposed De-LightSAM significantly surpasses EfficientViT-SAM [24] with a 56.47% and 25.92% Dice increase, respectively. On HD metric, our framework reduces the distance by up to 5.8$\times$ compared to current SOTA models, revealing more accurate boundary prediction. These results reveal the superior universal capabilities of our De-LightSAM on diverse medical segmentation tasks without the demand for manual prompts.
IV-C Comparison with State-of-the-arts on Unseen Domains